OpenStudy (ashtorah13):

A basketball player averaged 20 points a game over the course of six games. His scores in five of those games were 23, 18, 16, 24, and 27. How many points did he score in the sixth game?

OpenStudy (anonymous):

can you show your approach to this problem ?

OpenStudy (misssmartiez):

Ok, so what do we need to establish? The mean? average? If so what?

OpenStudy (ashtorah13):

I had no idea how to solve this problem. I actually left it blank. I need to find his point for the sixth game. At first I thought maybe making all of them the same number and what was left over would be able to work into something. In school we never did problems like this. Just finding averages.

OpenStudy (misssmartiez):

Hmm it seems to give the average though, are you sure that is what we need to find? "A basketball player averaged 20 points a game over the course of six games."

OpenStudy (misssmartiez):

We need to find the SIXTH number of the game. Do you know how ;)

OpenStudy (ashtorah13):

I am missing the sixth game point, I have to find what it is.

OpenStudy (lerenstudy):

So you have all the numbers up until the fifteenth game. If you add those up and then divide by 6, you'll get 18, right?

OpenStudy (anonymous):

Lets assume the sixth game point is "x" its the same problem as getting the average... \[\frac{ 23+18+16+24+27+x }{ 6} = 20\] find x :)

OpenStudy (misssmartiez):

1: figure out what we need to conduct to get the average, 20.

OpenStudy (lerenstudy):

You need to find the number that will bring that up to 20

OpenStudy (anonymous):

\[ \frac{23+ 18+ 16+ 24+ 27+x}{6} = 20 \]

OpenStudy (lerenstudy):

20 multiplied by six would get you 120, so the sum of the 6 games has to be 120

OpenStudy (misssmartiez):

As to what I was saying,

OpenStudy (ashtorah13):

haha after seeing how easy this was to set up I feel really dumb haha

OpenStudy (misssmartiez):

lol, anyway ;) 108 + _ _ / 6 = 20.

OpenStudy (misssmartiez):

Lets do this... 108 / 6 = 18. Now, what number added on 108, equals 20?

OpenStudy (ashtorah13):

12 Sorry for the slow response, was figuring out the problem.

OpenStudy (misssmartiez):

^ No problem!

OpenStudy (misssmartiez):

Correct!

OpenStudy (ashtorah13):

Thanks for all your help. Setting up algebra is terrible for me.

OpenStudy (misssmartiez):

No problem!